OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd.

OpenStudy (anonymous):

\[f(x) = 14^{3}\sqrt{x}\]

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

A function is even if f(-x) = f(x). A function is odd if f(-x) = -f(x). Substitute -x into f(x) and simplify. What do you get?